# Vulnerability Coordination Under the Cyber Resilience Act

### v2

@@ Abstract
A new Cyber Resilience Act (CRA) was recently agreed upon in the European Union (EU). It imposes many new cyber security requirements practically to all information technology products, whether hardware or software. The paper examines and elaborates the CRA’s new requirements for vulnerability coordination, including vulnerability disclosure. Although these requirements are only a part of the CRA’s obligations for vendors, also some new vulnerability coordination mandates are present, including particularly with respect to so-called actively exploited vulnerabilities. The CRA further alters the coordination practices on the side of public administrations. With the examination, elaboration, and associated discussion, the paper contributes to the study of cyber security regulations, providing also a few practical takeaways.
Keywords:
c
yber security regulations, vulnerability disclosure, vulnerability databases, actively exploited vulnerabilities, near misses, EU, CRA, NIS2
@@ 1 Introduction
The EU’s new Cyber Resilience Act, that is, Regulation (EU) 2024/2847, was agreed upon in October 2024. The paper presents an analysis of the CRA’s implications for vulnerability coordination, including vulnerability disclosure. Although the handling of vulnerabilities is a classical topic, the CRA imposes some new requirements. It also enlarges the scope of theoretical concepts involved. For these reasons, the paper carries both practical relevance and timeliness. In terms of explicitly related work, the paper continues the initial assessments [7] that were conducted based on the European Commission’s proposal for the CRA.
@@
The CRA is a part of a larger regulatory package for cyber security that was pursued by the von der Leyen’s first Commission, which was in office between 2019 and 2024. Although many of the other laws in the package are on the side of critical infrastructure, the overall goal of improved coordination, which can be understood broadly as a management of dependencies [17, 26], is strongly present throughout the package. Whether it is risk analysis or incident management, the EU context generally implies that some coordination is required at multiple levels; between cyber security professionals, companies, industry sectors, national public authorities, and EU-level institutions, among other abstraction levels. To make sense of the coordination involved, it is useful to analytically separate horizontal coordination (such as between the member states) and vertical coordination, which occurs between EU-level institutions and units at a national level [22, 23]. Then, as will be elaborated, the CRA involves both horizontal and vertical coordination. Coordination requires also technical infrastructures.
@@
Regarding the larger regulatory package for cyber security, relevant to mention is Directive (EU) 2022/2555, which is commonly known as the NIS2 directive. With respect to vulnerability coordination—a term soon elaborated in Section 2, the directive strengthens and harmonizes the role of national computer security incident response teams (CSIRTs) in Europe. In general, various networks between CSIRTs, both in Europe and globally, have been important for improving cyber security throughout the world, often overcoming obstacles that more politically oriented cyber security actors have had [28]. Importantly for the present purposes, the NIS2 directive, in its Article 12(2), specifies also an establishment of a new European vulnerability database operated by the European Union Agency for Cybersecurity (ENISA), which is the principal EU-level institution coordinating particularly with the national CSIRTs but also involved in international collaboration. While it is still too early to say with confidence, together the CRA and NIS2 may thus also affect the infrastructures for vulnerability archiving, which is currently strongly built upon institutions and companies in the United States, including particularly the not-for-profit MITRE corporation involved in the assignment of Common Vulnerabilities and Exposures (CVEs) and the National Vulnerability Database (NVD) maintained by the National Institute of Standards and Technology (NIST). With respect to the regulatory cyber security package, it can be further noted that the CRA is rather unique in a sense that its legal background and logic are motivated by the EU’s product safety laws and the general consumer protection jurisprudence [7]. Therefore, also the associated terminology is somewhat different compared to the other laws in the package. For instance, the CRA brings market surveillance authorities to the public administration of cyber security in the EU. In general, these authorities are responsible for ensuring that only compliant products enter into and circulate on the EU’s internal market; the examples include authorities ensuring the safety of food, drugs, aviation, chemicals, and so forth and so on.
@@
The CRA covers the whole information technology sector and beyond. Both hardware and software products with a networking functionality are in its scope. Among the notable exemptions not covered by the regulation, as specified in the CRA’s Article 2, are medical devices, motor vehicles, and ships and marine equipment. Furthermore, recital 12 clarifies that cloud computing is also excluded. The CRA categorizes products into three categories: “normal”, “important”, and “critical” products. While all products need to comply with the CRA’s essential requirements, more obligations are placed upon the important and critical products [24]. However, in what follows, the focus is only on the CRA’s implications for vulnerability coordination, including vulnerability disclosure. With this focus in mind, the opening Section 2 introduces the vulnerability coordination and disclosure concepts. Then, the subsequent Section 3 discusses and elaborates the CRA’s new implications to vulnerability coordination and disclosure. The implications are not uniform across everyone; therefore, Section 4 continues by briefly elaborating the requirements for open source software products and their projects. To some extent, the CRA addresses also the pressing issues with supply chain security. The important supply chain topic is discussed in Section 5. A conclusion and a discussion are presented in the final Section 6.
@@ 2 Vulnerability Disclosure and Coordination
There are various ways to disclose and coordinate a vulnerability between the vulnerability’s discoverer and a software vendor affected by the vulnerability. Historically—and perhaps still so—a common practice was so-called direct disclosure through which a discoverer and a vendor coordinate and negotiate privately, possibly without notifying any further parties, including maintainers of vulnerability databases and public authorities [25]. Partially due to the problems associated with direct disclosure, another historical—and sometimes still practiced—process was a so-called full disclosure via which a discoverer makes his or her vulnerability discovery public, possibly even without notifying a vendor affected or anyone else beforehand. One of the primary reasons for the full disclosure practice was—and sometimes still is—a deterrence against vendors who would not otherwise cooperate. However, this practice can be seen to increase security risks because sensitive information is publicly available but no one has had time to react; there are no patches available because a vendor may not even know about a vulnerability information being available in the public.
@@
With respect to Europe and the public sector, CSIRT-based coordinated vulnerability disclosure model, the NIS2 directive specified in its Article 12(1) three primary tasks for the national CSIRTs in the member states: (1) they should identify and contact all entities concerned, including the vendors affected in particular; (2) they should assist natural or legal persons who report vulnerabilities; and (3) they should negotiate disclosure timelines and arrange coordination for vulnerabilities that affect multiple entities. While simple, these legal mandates for public sector bodies are important because the legal landscape for vulnerability disclosure has been highly fragmented in Europe [5, 10]. The CRA relies partially upon these mandates in NIS2, but it also imposes new requirements for both national CSIRTs and vendors (or manufacturers, as they are called in the CRA). Before continuing, it should be clarified that a broader term of vulnerability coordination has sometimes been preferred in the literature because vulnerability disclosure presents only a narrow perspective on the overall handling of vulnerabilities [26]. In addition to contacting, communicating, assisting, and negotiating, numerous other tasks are typically involved, including development of patches, testing these, and delivering these to customers, clients, and users in general, assignment of CVE identifiers, writing and releasing security advisories, archiving information to vulnerability databases, and so forth and so on. These additional tasks do not affect only vendors but also CSIRTs often need to carry out additional tasks beyond those specified in the NIS2 directive. For these reasons, also the present work uses a term vulnerability coordination.
3 The CRA’s New Requirements
@@
The CRA separates voluntary vulnerability disclosure from mandatory obligations placed upon vendors. Regardless whether a disclosure of a vulnerability is done on voluntary or mandatory basis, it should be primarily done through a national CSIRT, although voluntary disclosure may be done also to ENISA directly according to the CRA’s Article 15(1). Then, according to Article 15(2), both vendors and others may also report not only vulnerabilities but also security incidents and so-called near misses. These near misses are defined in the NIS2’s Article 6(5) as events that “could have compromised the availability, authenticity, integrity or confidentiality of stored, transmitted or processed data or of the services offered by, or accessible via, network and information systems” but which were successfully mitigated. These mitigated security events are one but not the only example of new terminology brought by the EU’s new laws.
@@
When a disclosure is done to a national CSIRT, the given CSIRT should then without undue delay carry out vertical coordination toward the EU-level. This important obligation placed upon national public sector authorities is specified in the CRA’s Article 16(1) according to which a common EU-level disclosure infrastructure is established. It is maintained by ENISA. Through this infrastructure, according to Article 16(2), a given national CSIRT should notify other European CSIRTs designed as public sector authorities on those territories on which a vendor’s vulnerable products have been made available. The territorial notion seems sensible enough when dealing with physical products, whether routers, mobile phones, or automobiles. However, many digital products, including software products in particular, are distributed online without any specific territorial scope. For such products, it seems a reasonable assumption that all European CSIRTs should be notified. This point applies particularly to large software vendors, such as, say, Microsoft or Google. Similar reasoning was behind a schism during the CRA’s political negotiations because some companies and member states feared that sensitive vulnerability information would unnecessarily spill over throughout Europe due to the EU-level reporting infrastructure [23]. The schism was resolved by relaxing the obligation placed upon CSIRTs. In particular, according to Article 16(2), national CSIRTs may delay coordination and reporting toward the EU-level under exceptional circumstances requested by a vendor affected. The article specifies three clarifying conditions regarding these exceptional circumstances: (1) a given vulnerability is actively exploited by a malicious actor and no other member state is affected; (2) immediate reporting would affect the essential interests of a member state; and (3) imminent high cyber security risk is stemming from reporting. Although a national CSIRT should always still notify ENISA about its decision to delay reporting, it seems that the three conditions are so loose that most cases could be delayed in principle. Time will tell how the regulators interpret the exceptional circumstances in practice.
@@
The exceptional circumstances noted contain the concept of actively exploited avulnerabilities. Such vulnerabilities are a core part of the CRA. An actively exploited vulnerability is defined in Article 3(42) as “a vulnerability for which there is reliable evidence that a malicious actor has exploited it in a system without permission of the system owner”. Recital 68 clarifies that security breaches, including data breaches affecting natural persons, are a typical example. The recital also notes that the concept excludes good faith testing, investigation, correction, or disclosure insofar as no malicious motives have been involved. Then, according to Article 14(1), all vendors should notify about actively exploited vulnerabilities in their products to both a given national CSIRT and ENISA. Article 14(2) continues by specifying that vendors should deliver an early notification within 24 hours, an update within 72 hours, and a final report within two weeks. Furthermore, the subsequent paragraphs 3 and 4 in Article 14 obligate vendors to report not only about actively exploited vulnerabilities but also about actual security incidents. These are defined in the NIS2’s Article 6(6) analogously to the near misses but without the successful mitigation condition. Furthermore, Article 14(8) specifies that both actively exploited vulnerabilities and security incidents should be communicated to the users impacted, possibly including the general public, either by a vendor affected or a given CSIRT.
@@
It is important to further note that the concept of actively exploited vulnerabilities is used also in the United States. In particular, the Cybersecurity & Infrastructure Agency (CISA) therein maintains a catalog of known exploited vulnerabilities [8]. Given the ENISA’s and CISA’s established cooperation [11], it remains to be seen whether the catalog will be used also for enforcing the CRA or otherwise helping its implementation. An analogous point applies to the international coordination between vulnerability databases, including particularly between the NVD and the EU’s upcoming vulnerability database enacted in the NIS2 directive. A related important point is that less strict vulnerability reporting requirements are imposed upon open source software projects and their supporting foundations, which are known as stewards in the CRA’s parlance.
@@
Finally, it is important to emphasize that the CRA’s so-called essential cyber security requirements, which, like with the EU’s product safety laws [22], will be augmented in the nearby future with technical standards, contain also obligations explicitly related to vulnerabilities and their coordination. In general, as specified in Annex I, vendors should only make available products without known exploitable vulnerabilities, but they should also ensure that vulnerabilities can be addressed through security updates, including, where applicable, automated security updates separated from functional updates that bring new features. They must also establish and enforce a coordinated vulnerability disclosure policy, and apply effective and regular security testing and reviews, among other things. As soon discussed in Section 5, the essential cyber security requirements imply also new requirements for supply chain management and coordination.
4 The CRA and Open Source

### v3

@@ Abstract
The Cyber Resilience Act (CRA) of the European Union (EU) imposes many new cyber security requirements practically to all network-enabled information technology products, whether hardware or software. The paper examines and elaborates the CRA’s new requirements for vulnerability coordination, including vulnerability disclosure. Although these requirements are only a part of the CRA’s obligations for vendors, also some new vulnerability coordination mandates are present. In particular, so-called actively exploited vulnerabilities require mandatory reporting. In addition to elaborating the reporting logic, the paper discusses the notion of actively exploited vulnerabilities in relation to the notion of known exploited vulnerabilities used in the United States. The CRA further alters the coordination practices on the side of public administrations. The paper addresses also these new practices. With the examination elaboration, and associated discussion based on conceptual analysis, the paper contributes to the study of cyber security regulations, providing also a few takeaways for further research.
Keywords: vulnerability disclosure, vulnerability databases, known vulnerabilities, actively exploited vulnerabilities, known exploited vulnerabilities
@@ 1 Introduction
The EU’s Cyber Resilience Act, that is, Regulation (EU) 2024/2847, was agreed upon in October 2024. The paper presents an analysis of the CRA’s implications for vulnerability coordination, including vulnerability disclosure. Although the handling of vulnerabilities is a classical research topic, the CRA imposes some new requirements that warrant attention. It also enlarges the scope of theoretical concepts involved. For these reasons, the paper carries both practical relevance and timeliness. Regarding timeliness, the reporting obligations upon severe incidents and actively exploited vulnerabilities will be enforced already from September 2026 onward. Regarding practical relevance, it is worth elaborating and clarifying the obligations and concepts involved because some vendors have had difficulties with them due to a lack of expertise and other reasons [25]. In terms of explicitly related work, the paper continues the initial assessments [7] that were conducted based on the European Commission’s proposal for the CRA. By using conceptual analysis [21], the research question examined is: what new requirements the CRA entails for vulnerability coordination? To the best of the authors’ knowledge, the question has not been previously asked and answered.
@@
The CRA is a part of a larger regulatory package for cyber security that was pursued by the von der Leyen’s first Commission, which was in office between 2019 and 2024. Although many of the other regulations and directives in the package are on the side of critical infrastructure, the overall goal of improved coordination, which can be understood broadly as a management of dependencies [18, 31], is strongly present throughout the package. Whether it is risk analysis or incident management, the EU context generally implies that some coordination is required at multiple levels; between cyber security professionals, companies, industry sectors, national public authorities, and EU-level institutions, among other abstraction levels [32]. To make sense of the coordination involved, it is useful to analytically separate horizontal coordination (such as between the member states) and vertical coordination, which occurs between EU-level institutions and units at a national level [26, 27]. Then, as will be elaborated, the CRA involves both horizontal and vertical coordination. Coordination requires also technical infrastructures.
@@
Regarding the larger regulatory package for cyber security, relevant to mention is Directive (EU) 2022/2555, which is commonly known as the second network and information systems (NIS2) security directive. With respect to vulnerability coordination—a term soon elaborated in Section 2, the directive strengthens and harmonizes the role of national computer security incident response teams (CSIRTs) in Europe. In general, various networks between CSIRTs, both in Europe and globally, have been important for improving cyber security throughout the world, often overcoming obstacles that more politically oriented cyber security actors have had [35]. Importantly for the present purposes, the NIS2 directive, in its Article 12(2), specifies also an establishment of a new European vulnerability database operated by the European Union Agency for Cybersecurity (ENISA), which is the principal EU-level institution coordinating particularly with the national CSIRTs but also involved in international collaboration. Although the database is already online [12], it is still too early to say with confidence whether the CRA and NIS2 may affect the infrastructures for vulnerability archiving, which is currently strongly built upon institutions and companies in the United States, including particularly the not-for-profit MITRE corporation involved in the assignment of Common Vulnerabilities and Exposures (CVEs) and the National Vulnerability Database (NVD) maintained by the National Institute of Standards and Technology (NIST) of the United States.
With respect to the EU’s regulatory cyber security package, it can be further noted that the CRA is rather unique in a sense that its legal background and logic are motivated by the EU’s product safety law and the general consumer protection jurisprudence [7]. Therefore, also the associated terminology is somewhat different compared to the other legislative acts in the package. For instance, the CRA brings market surveillance authorities to the public administration of cyber security in the EU. In general, these authorities are responsible for ensuring that only compliant products enter into and circulate on the EU’s internal market; the examples include authorities ensuring the safety of food, drugs, aviation, chemicals, and so forth.
@@
The CRA covers the whole information technology sector and beyond. Both hardware and software products with a networking functionality are in its scope. Among the notable exemptions not covered by the regulation, as specified in the CRA’s Article 2, are medical devices, motor vehicles, and ships and marine equipment. Furthermore, recital 12 clarifies that cloud computing is also excluded. The CRA categorizes products into three categories: “normal”, “important”, and “critical” products. While all products need to comply with the CRA’s essential requirements, which have been analyzed in recent research [30, 28], more obligations are placed upon the important and critical products  However, in what follows, the focus is only on the CRA’s implications for vulnerability coordination, including vulnerability disclosure. With this focus in mind, the opening Section 2 introduces the vulnerability coordination and disclosure concepts. Then, the subsequent Section 3 discusses and elaborates the CRA’s new implications to vulnerability coordination and disclosure. The implications are not uniform across everyone; therefore, Section 4 continues by briefly elaborating the requirements for open source software products and their projects. To some extent, the CRA addresses also the pressing issues with supply chain security. The important supply chain topic is discussed in Section 5. A conclusion is presented in Section 6. The final Section 7 discusses further research.
@@ 2 Vulnerability Disclosure and Coordination
There are various ways to disclose and coordinate a vulnerability between the vulnerability’s discoverer and a software vendor affected by the vulnerability. Historically—and perhaps still so—a common practice was so-called direct disclosure through which a discoverer and a vendor coordinate and negotiate privately, possibly without notifying any further parties, including maintainers of vulnerability databases and public authorities [29]. Partially due to the problems associated with direct disclosure, another historical—and sometimes still practiced—process was a so-called full disclosure via which a discoverer makes his or her vulnerability discovery public, possibly even without notifying a vendor affected or anyone else beforehand. One of the primary reasons for the full disclosure practice was—and sometimes still is—a deterrence against vendors who would not otherwise cooperate [29, 34]. However, this practice can be seen to increase security risks because sensitive information is publicly available but no one has had time to react; there are no patches available because a vendor may not even know about a vulnerability information being available in the public.
@@
With respect to Europe and the CSIRT-based coordinated vulnerability disclosure model, the NIS2 directive specifies in its Article 12(1) three primary tasks for the national CSIRTs in the member states: (1) they should identify and contact all entities concerned, including the vendors affected in particular; (2) they should assist natural or legal persons who report vulnerabilities; and (3) they should negotiate disclosure timelines and arrange coordination for vulnerabilities that affect multiple entities. While simple, these legal mandates for public sector bodies are important because the legal landscape for vulnerability disclosure has been highly fragmented in Europe [5, 10]. The CRA relies partially upon these mandates in NIS2, but it also imposes new requirements for both national CSIRTs and vendors.
Before continuing further, it should be clarified that a broader term of vulnerability coordination has sometimes been preferred in the literature because vulnerability disclosure presents only a narrow perspective on the overall handling of vulnerabilities [31]. In addition to contacting, communicating, assisting, and negotiating, numerous other tasks are typically involved, including development of patches, testing these, and delivering these to customers, clients, and users in general, assignment of CVE identifiers, writing and releasing security advisories, archiving information to vulnerability databases, and so forth and so on. These additional tasks do not affect only vendors; also CSIRTs may need to carry out additional tasks beyond those specified in NIS2. For these reasons, also the present work uses the vulnerability coordination term.
3 The CRA’s Vulnerability Coordination Requirements
@@
The CRA separates voluntary vulnerability disclosure from mandatory obligations placed upon vendors. Regardless whether a disclosure of a vulnerability is done on voluntary or mandatory basis, it should be primarily done through a national CSIRT, although voluntary disclosure may be done also to ENISA directly according to the CRA’s Article 15(1). Then, according to Article 15(2), both vendors and others may also report not only vulnerabilities but also severe security incidents and so-called near misses. These near misses are defined in the NIS2’s Article 6(5) as events that “could have compromised the availability, authenticity, integrity or confidentiality of stored, transmitted or processed data or of the services offered by, or accessible via, network and information systems” but which were successfully mitigated. These mitigated security events are one but not the only example of new terminology brought by the EU’s new cyber security law.
When a disclosure is done to a national CSIRT, the given CSIRT should then without undue delay carry out vertical coordination toward the EU-level. This important obligation placed upon national public sector authorities is specified in the CRA’s Article 16(1) according to which a common EU-level disclosure infrastructure is established. It is maintained by ENISA. Through this infrastructure, according to Article 16(2), a given national CSIRT should notify other European CSIRTs designed as public sector authorities on those territories on which a vendor’s vulnerable products have been made available.
@@
The territorial notion seems sensible enough when dealing with physical products, whether routers, mobile phones, or automobiles. However, many digital products, including software products in particular, are distributed online without any specific territorial scope. For such products, it seems a reasonable assumption that all European CSIRTs should be notified. This point applies particularly to large software vendors. Similar reasoning was behind a schism during the CRA’s political negotiations because some companies and member states feared that sensitive vulnerability information would unnecessarily spill over throughout Europe due to the EU-level reporting infrastructure [27]. The schism was resolved by relaxing the obligation placed upon CSIRTs.
In particular, according to Article 16(2), national CSIRTs may delay coordination and reporting toward the EU-level under exceptional circumstances requested by a vendor affected. The article specifies three clarifying conditions regarding these exceptional circumstances: (1) a given vulnerability is actively exploited by a malicious actor and no other member state is affected; (2) immediate reporting would affect the essential interests of a member state; and (3) imminent high cyber security risk is stemming from reporting. Although a national CSIRT should always still notify ENISA about its decision to delay reporting, it seems that the three conditions are so loose that most cases could be delayed in principle. Time will tell how the regulators interpret the exceptional circumstances in practice.
@@
The exceptional circumstances noted contain the concept of actively exploited vulnerabilities. Such vulnerabilities are a core part of the CRA. An actively exploited vulnerability is defined in Article 3(42) as “a vulnerability for which there is reliable evidence that a malicious actor has exploited it in a system without permission of the system owner”. Recital 68 clarifies that security breaches, including data breaches affecting natural persons, are a typical example. The recital also notes that the concept excludes good faith testing, investigation, correction, or disclosure insofar as no malicious motives have been involved.
Then, according to Article 14(1), all vendors should notify about actively exploited vulnerabilities in their products to both a given national CSIRT and ENISA. Article 14(2) continues by specifying that vendors should deliver an early notification within 24 hours, an update within 72 hours, and a final report within two weeks. As could be expected, concerns have been expressed among vendors about these strict deadlines [25]. Also reporting quality remains a question mark [32]. In any case, the subsequent paragraphs 3 and 4 in Article 14 obligate vendors to report not only about actively exploited vulnerabilities but also about severe security incidents. These are defined in the NIS2’s Article 6(6) analogously to the near misses but without the successful mitigation condition. Furthermore, Article 14(8) specifies that both actively exploited vulnerabilities and severe security incidents should be communicated to the users impacted, possibly including the general public, either by a vendor affected or a given CSIRT.
@@
It is important to further note that the concept of actively exploited vulnerabilities is used implicitly also in the United States. In particular, the Cybersecurity & Infrastructure Agency (CISA) therein maintains a catalog of known exploited vulnerabilities (KEVs) [8]. Given the ENISA’s and CISA’s established cooperation [11], it remains to be seen whether the catalog will be used also for enforcing the CRA or otherwise helping its implementation. An analogous point applies to the international coordination between vulnerability databases, including particularly between the NVD and the EU’s new vulnerability database. A related important point is that less strict vulnerability reporting requirements are imposed upon open source software projects and their supporting foundations, which are known as stewards in the CRA’s parlance.
Regarding KEVs, the available documentation [8] says that three conditions must be satisfied for a vulnerability to be a KEV. First, a KEV must have a CVE identifier assigned. Then, exploitability is understood along the lines of the Common Vulnerability Scoring System (CVSS) as the ease of an attacker to take advantage of a CVE-referenced vulnerability. Thus, exploitability varies according to a number of factors, such as whether a vulnerability can be exploited remotely through a network, whether a proof-of-concept exploit is available, whether authentication or privileges are required for exploitation, and so forth and so on. A vulnerability that cannot be exploited under reasonable constraints cannot be a KEV. Second, however, for an exploitable vulnerability to be a KEV, it must have been either exploited or under active exploitation Both terms require that there is “reliable evidence that execution of malicious code” has occurred [8]. Such evidence does not necessarily mean that a vulnerability has been successfully exploited; a KEV can refer also to an attempted exploitation, including those conducted in a honeypot or related deception environment. It can be also noted that execution of malicious code rules out network scanning and denial of service attacks. Third, a KEV must have clear remediation guidelines for it to be accepted into the catalog. These range from updates provided by vendors to temporary mitigations and workarounds. It will be interesting to see whether also the CRA’s concept of actively exploited vulnerabilities will be tied to these or closely related principles or whether a new European conceptual framework will be developed in the future.
@@
Finally, it is important to emphasize that the CRA’s so-called essential cyber security requirements, which, like with the EU’s product safety law [26], will be accompanied with harmonized standards, contain also obligations explicitly related to vulnerabilities and their coordination. In general, as specified in Annex I, vendors should only make available products without known exploitable vulnerabilities, but they should also ensure that vulnerabilities can be addressed through security updates, including, where applicable, automated security updates separated from functional updates that bring new features [28]. They must also establish and enforce a coordinated vulnerability disclosure policy, and apply effective and regular security testing and reviews, among other things. As soon discussed in Section 5, the essential cyber security requirements imply also new requirements for supply chain management.
4 The CRA’s Vulnerability Coordination Requirements for Open Source Projects